November 10th photos resend.
Crew 239 Journalist Report for November 10th
Journalist Report 10 Nov 2021
By Emily Milne
MDRS Crew 239 Sol 3 (11/10/2021):
We began our morning with a breakfast of oatmeal, Cliff bars, tea, and coffee before our 1000 EVA. The EVA was a return to rover training by Kissing Camel Ridge and consisted of team leader GHO Cynthia Montanez, HSO Emily Milne, and ENG Aidan Moga. ENG Moga took samples for analysis in his spectrometer to find potential landing sites. Upon returning at 1202 CO Crisman made jambalaya for Lunch, GHO Montanez transplanted and watered the plants, ENG Moga utilized the spectrometer, and the rest of the crew worked on individual research. HSO Milne cooked pasta with gravy for dinner using tomato powder, tomato chunks, onion, and pork sausage. After eating, the crew watched “The Thing” before retiring for the night.
Crew 239 Operations Report for November 10th
Crew 239 Operations Report 10-11-2021
Name of person filing report: Aidan Moga
Non-nominal systems: Toilet, Propane supply to generator, water heater
Action taken for non-nominal systems: The leak in the toilet fitting has gotten worse. The crew is working on better diagnosing and solving the problem later tonight. The power supply situation was better today, due to the sunny weather. The crew was able to charge the rovers and devices during the daytime. Power rationing continues in the afternoon/evenings. At around 18:20 the water heater was on and started to leak gas into the lower level of the hab. The crew attempted to use the gas detector from the ram to locate the leak. However, the leak couldn’t be replicated and wasn’t precisely located. The crew will continue to monitor the situation.
Daily crew water use: 21 gallons
GreenHab water use: 3 gallons
Hab toilet tank emptied: No
Rovers used: Yes, Spirit and Opportunity.
Notes on rovers (report nominal or any anomalies): Nominal
Summary of crew and Mission Support internet: Nominal
Summary of suits and radios: During EVA pause, the fan on GHO’s suit 7 sounded like it was being interrupted. The suit was set aside and replaced with suit 8 for the EVA. When suit 7 was tested later, the fan appeared to operate nominally. For now, suit 8 will continue to be used. Some of the radios have damaged antennas; these have been put aside to not be used during EVAs.
Campus wide inspection, if action taken, what and why? Nominal
Summary of general operations: Nominal
Summary of Hab operations: Nominal
Summary of GreenHab operations (note any harvested crops):
· Temperature throughout Sol 2:
o 1:00 p.m. -> 110 degree F with 7% humidity
o 4:34 p.m. -> 72.5 degree F with 19% humidity
o 6:30 p.m. -> 70.6 degree F with 19 % humidity
· Hours of Supplemental Light: None
· Time(s) of watering for crops:
o At 1:19 p.m. I used 1 gallon
o At 4:54 p.m. I used 3 gallons
o Transplanted Thai pepper plant to a larger pot
o Transplanted largest cucumber plant to a larger pot and straightened the stems.
o Moved smaller cucumbers to a location furthest away from the door to help with climate control.
o Transplanted two large sweetie tomatoes into seven smaller pots
o Transplanted regular, but undefined, tomato plant to a larger pot.
o Discovered that the zucchini’s do not have a disease, instead they are ‘burning’ due to the extreme temperatures the GreenHab has. I will handle accordingly to prevent any more damage on these plants.
· Support/ Supplied Needed:
o Requesting more potting soil
o Requesting red wagon within the GreenHab to be removed.
Summary of ScienceDome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns, supplies needed and requests: None
Crew Photos for November 10th
Greetings from Mars.
Here are the pictures from Sol 3. Our Picture of the Day is "Sunrise".
Brian Murphy, Crew Artis